Fermi National Laboratory


Accelerator Update for May 11, 2005 - May 13, 2005

Wednesday May 11
The day shift began with Operations monitoring store 4136, with the Antiproton Source (Pbar) stacking, with the Recycler in shot setup, and with MiniBooNE taking beam.

The Recycler finished stashing antiprotons and Pbar resumed stacking and NuMI resumed taking beam at 10:34 AM.

At 10:54 AM, a Booster power supply (EAPS) tripped off for the first of two times in five minutes. It reset okay both times.

At 11:05 AM, a Booster RF station (BRF4) tripped off. When operators reset the station, the EAPS tripped off. An expert investigated.

Pbar resumed stacking at 11:40 AM. Booster experts reported that they brought up BRF18 and took down BRF4.

At 1:40 PM, BRF8 tripped off for the first of many trips over the next twenty-six minutes. On the last trip at 2:06 PM RF stations 13, 15, and 16 tripped off too. Experts investigated and everything reset okay.

Pbar experts began a study period at 1:44 PM. Operators disabled stacking for the study.

At 1:51 PM, NuMI tripped off due to oscillations in the target and baffle LCW systems. Experts investigated.

At 2:11 PM, Operations reported that Booster was down due to BRF8 tripping off. BRF8 had tripped off due to a vacuum indication, but operators found no vacuum problem. Experts continued their investigations.

NuMI resumed taking beam at 4:31 PM.

A Main Injector (MI) expert began a study at 2:55 PM.

Pbar resumed stacking at 3:19 PM.

At 4:54 PM, BRF5 tripped off and two minutes later BRF3 tripped off. They both reset okay.

MI studies finished at 7:37 PM.

Recycler began stashing antiprotons at 10:04 PM and completed the stash thirty minutes later. Pbar resumed stacking and NuMI and MCenter resumed taking beam at 10:47 PM.

Thursday May 12
The midnight shift began with Operations monitoring store 4136 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations terminated store 4136 at 4:02 AM.

Tevatron (TeV) experts began a chromaticity study at 4:25 AM.

At 4:49 AM, BRF16 tripped off for the first of two times in four minutes. It reset okay both times.

At 5:21 AM, the MI-8 transfer line tripped off due to equipment failure. This halted all downstream beam. Operators replaced the equipment and stacking and beam to MiniBooNE, NuMI, and MCenter resumed by 5:48 AM.

Operations disabled the MiniBooNE beam at 6:08 AM for a scheduled access.

TeV experts completed their studies at 8:36 AM.

Operators prepared the Collider Detector at Fermilab (CDF) and D-Zero (D0) collision halls for access at 8:52 AM.

Both CDF and D0 were out of access by 1:31 PM.

Operations put the TeV through a dry squeeze at 2:04 PM.

Booster experts reported at 2:40 PM that BRF4 was back up and online.

Operations put the TeV into shot setup around 4 PM.

At 4:52 PM, operators turned off the Booster beam to access the tunnel and replace some equipment.

Operations established store 4138 at 5:16 PM with an initial luminosity of 126.97E30. This store set a NEW RECORD LUMINOSITY.

Operators came out of Booster and stacking resumed at 5:47 PM. Beam to NuMI resumed a few minutes later.

Operations established beam to MCenter at 7:32 PM.

A Pbar expert began Debuncher orbit studies at 10:36 PM.

Operations put the Recycler into shot setup to stash antiprotons at 11:09 PM.

Friday May 13
The midnight shift began with Operations monitoring store 4138, with the Recycler stashing, and with MiniBooNE taking beam.

The recycler stash ended and Pbar resumed stacking and NuMI resumed taking beam at 12:05 AM.

MCenter resumed taking beam at 12:12 AM.

At 1:53 AM, a thunderstorm rolled across Fermilab. The TeV suffered a sector F4 quench due to a lightening strike. A few minutes later we lost the MI power supply beam permit due to another strike and worse, we also lost the stack.

At 1:56 AM, Operations reported several power glitches. From the Linac to the TeV, Operations reported that many systems had tripped off.

At 2:22 AM, Operations reported that they couldn't get BRF18 to reset.

Operations established beam to MiniBooNE at 2:57 AM.

Operations established beam to NuMI at 3:17 AM and to MCenter at 3:20 AM.

Operations put the TeV through an extended dry squeeze at 4:40 AM.

Pbar resumed stacking at 6:12 AM.

At 6:47 AM, operators halted MiniBooNE beam at their request. They think there might be a problem with their horn's LCW system. Experts were called in.

TeV experts began a study period at 7:33 AM.

The Plan for Friday and the Weekend
The plan for today is to stack and stash until late this afternoon or early this evening and then establish the next store. The weekend will consist of stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/13/2005   email Fermilab